For the cost function C(Q) = 75 + 4Q + 2Q2, the marginal cost of producing 5 units of output is:
Answer #1

Variable cost = 4Q + 2Q^2

Marginal cost is the differentiation of variable cost with respect to Q

Marginal cost = 4 + 4Q

When Q= 5

Marginal cost = 4 + 4(5)= 24